# Untangling Lariats: Subgradient Following of Variationally Penalized Objectives

### v1

@@ Appendix B Dual Proximal Gradient
We give here further details on the dual proximal gradient method (DPG) we used in comparing our core algorithmic contribution. DPG merits further investigation on its own and to the best of our knowledge the short analysis with general norm penalties below was not derived in previous research.
@@
$\min_{{\mathbf{x}}_{1:n}}~{}\frac{1}{2}\sum_{i=1}^{n}\|{\mathbf{x}}_{i}-\mathbf{y}_{i}\|_{2}^{2}+\sum_{i=1}^{n-1}\|{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\|~{}.$
@@
$\max_{\boldsymbol{\alpha}_{0:n}}~{}\sum_{i=1}^{n}\|\mathbf{y}_{i}\|_{2}^{2}-\|\mathbf{y}_{i}+\boldsymbol{\alpha}_{i}-\boldsymbol{\alpha}_{i-1}\|_{2}^{2}~{}\;\text{s.t.}\;~{}\|\boldsymbol{\alpha}_{i}\|_{*}\leq 1~{}.$
For boundary conditions we set $\boldsymbol{\alpha}_{0}\equiv\boldsymbol{\alpha}_{n}\equiv\mathbf{0}$. Denote by $B_{*}=\{{\mathbf{x}}:\|{\mathbf{x}}\|_{*}\leq 1\}$ the unit ball w.r.t the dual norm and $\Pi_{B_{*}}(\cdot)$ be the projection operator onto the unit ball. In each iteration of dual proximal gradient,
$\boldsymbol{\alpha}_{i}\,\leftarrow\,\Pi_{B_{*}}\big{(}\boldsymbol{\alpha}_{i}-\eta(\boldsymbol{\alpha}_{i}-\boldsymbol{\alpha}_{i-1}+\mathbf{y}_{i}+\boldsymbol{\alpha}_{i}-\boldsymbol{\alpha}_{i+1}-\mathbf{y}_{i+1})\big{)}~{}.$
In order to underscore the advantage of the sub-gradient following method, let us briefly examine the condition number of the dual problem. Denote by $\mathsf{D}$ the difference matrix
$\mathsf{D}=\begin{bmatrix}-1&1&0&\dots&0&0\\ 0&-1&1&\dots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\dots&-1&1\end{bmatrix}\in\mathsf{R}^{n-1\,\times\,n}~{}.$
@@
$\alpha_{i}=\frac{{(-1)}^{i}}{\sqrt{n-1}}~{},$
we have $\boldsymbol{\alpha}^{\top}\mathsf{D}\mathsf{D}^{\top}\boldsymbol{\alpha}=4-\frac{2}{n-1}$. Assume without loss of generality that $n$ is even and let us choose next
$\alpha_{i}=\frac{\frac{n}{2}-|i-\frac{n}{2}|}{\sqrt{n^{3}/12}}~{},$
for which we have $\|\boldsymbol{\alpha}\|_{2}\geq 1$ and $\boldsymbol{\alpha}^{\top}\mathsf{D}\mathsf{D}^{\top}\boldsymbol{\alpha}=12/n^{2}$. Therefore, for $n\geq 3$ the largest Eigen value of $\mathsf{D}\mathsf{D}^{\top}$ is at least $3$ whereas the smallest Eigen value is at most $\mathcal{O}(n^{-2})$. Thus, the condition number of the dual problem is $\Omega(n^{2})$ which renders dual ascent algorithms, and in particular, DPG slow to converge.
@@ Appendix C Upper Bounds on Slack Variables
Denote the optimum of the original problem by ${\mathbf{x}}_{1:n}^{*}$, we claim that $\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}\leq\|\mathbf{y}_{i+1}-\mathbf{y}_{i}\|_{\infty}$ and thus we can set $D_{i}=\|\mathbf{y}_{i+1}-\mathbf{y}_{i}\|_{\infty}$. If $\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}=0$, the bound clearly holds. Otherwise let $S=\{j:|x_{i+1,j}^{*}-x_{i,j}^{*}|=\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}\}$. Consider $j\in S$ and let $s_{i,j}=\text{sign}(x_{i+1,j}^{*}-x_{i,j}^{*})$. From the optimality condition, there exists $\boldsymbol{\alpha}_{1:n-1}$ such that
@@
$=$ $y_{i+1,j}-y_{i,j}+\alpha_{i+1,j}-2\alpha_{i,j}+\alpha_{i-1,j}~{}.$
@@
$\leq$ $|y_{i+1,j}-y_{i,j}|+|\alpha_{i+1,j}|+|\alpha_{i-1,j}|-2|\alpha_{i,j}|~{}.$
@@
$\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}\;\leq\;\frac{1}{|S|}\big{(}\sum_{j\in S}|y_{i+1,j}-y_{i,j}|+\|\boldsymbol{\alpha}_{i+1}\|_{1}+\|\boldsymbol{\alpha}_{i-1}\|_{1}-2\big{)}\;\leq\;\|\mathbf{y}_{i+1}-\mathbf{y}_{i}\|_{\infty}~{},$
using the fact that for $j\notin S:\alpha_{i,j}=0$ and $\|\boldsymbol{\alpha}_{i}\|_{1}=1$ since $\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}\neq 0$.
@@ Appendix D Iterative Algorithm for High-Order Variational Penalties
$\frac{1}{2}\|{\mathbf{x}}-\mathbf{y}\|^{2}+\sum_{i=1}^{n-2}\lambda_{i}|(x_{i+2}-x_{i+1})-(x_{i+1}-x_{i})|^{2}~{}~{}.$
@@
${\mathbf{x}}^{t+1}=\displaystyle\operatorname*{\arg\!\min}_{\mathbf{x}}\frac{1}{2}\|{\mathbf{x}}-\mathbf{y}\|^{2}+\sum_{i=1}^{n-2}\frac{|(x_{i+2}-x_{i+1})-(x_{i+1}-x_{i})|^{2}}{2\rho_{i}^{t}}$
where $\rho_{i}^{t}=\displaystyle\max\big{(}|(x_{i+2}^{t}-x_{i+1}^{t})-(x_{i+1}^{t}-x_{i}^{t})|,\,\epsilon\big{)}$.
@@
$|(x_{i+2}-2x_{i+1}+x_{i})-(x^{\prime}_{i+2}-2x^{\prime}_{i+1}+x^{\prime}_{i})|^{2}\leq 4\left(|x_{i+2}-x^{\prime}_{i+2}|^{2}+2|x_{i+1}-x^{\prime}_{i+1}|^{2}+|x_{i}-x^{\prime}_{i}|^{2}\right)~{}.$
@@
$\tilde{\mathsf{D}}=\begin{bmatrix}-1&1&0&\dots&0&0\\ 0&-1&1&\dots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\dots&-1&1\\ 0&0&0&\dots&0&0\\ \end{bmatrix}\in\mathsf{R}^{n\,\times\,n}~{}.$
@@
$\operatorname*{\arg\!\min}_{\mathbf{x}}\frac{1}{2}\|{\mathbf{x}}-\mathbf{y}\|^{2}+\big{\|}\big{[}\mathsf{\tilde{D}}^{k}{\mathbf{x}}\big{]}_{1:n-k}\big{\|}_{1}~{}.$ (17)
In terms of run time, we “pay” a factor of $4^{k}$ for the number of iterations required to converge and each iteration takes $\mathcal{O}(k^{2}n)$ for solving the linear system. In comparison, when using the dual proximal gradient method, the complexity of each iteration grows as $\mathcal{O}(\gamma^{k})$ where $\gamma=\Omega(n^{2})$ is the condition number of the first order problem. Pseudocode for $k>0$ is provided below.
Algorithm 2 $k^{\text{th}}$-Order Nariat

### v2

@@ Appendix B Dual Proximal Gradient
We give here further details on the dual proximal gradient method (DPG) used in the empirical studies. DPG merits further investigation on its own and to the best of our knowledge the short analysis with general norm penalties below was not derived in previous research.
@@
$\min_{{\mathbf{x}}_{1:n}}~{}\frac{1}{2}\sum_{i=1}^{n}\big{\|}{\mathbf{x}}_{i}-\mathbf{y}_{i}\big{\|}_{2}^{2}\;+\;\sum_{i=1}^{n-1}\big{\|}{\mathbf{x}}_{i+1}-{\mathbf{x}}_{i}\big{\|}~{}~{}.$
@@
$\max_{\boldsymbol{\alpha}_{0:n}}~{}\sum_{i=1}^{n}\big{\|}\mathbf{y}_{i}\big{\|}_{2}^{2}-\big{\|}\mathbf{y}_{i}+\boldsymbol{\alpha}_{i}-\boldsymbol{\alpha}_{i-1}\big{\|}_{2}^{2}~{}\;\text{s.t.}\;~{}\big{\|}\boldsymbol{\alpha}_{i}\big{\|}_{*}\leq 1~{}~{}.$
For boundary conditions we set $\boldsymbol{\alpha}_{0}\equiv\boldsymbol{\alpha}_{n}\equiv\mathbf{0}$. Denote by $B_{*}=\{{\mathbf{x}}:\|{\mathbf{x}}\|_{*}\leq 1\}$ the unit ball w.r.t the dual norm. Let $\Pi_{B_{*}}(\cdot)$ denote the projection operator onto the unit ball. In each iteration of dual proximal gradient,
$\boldsymbol{\alpha}_{i}\,\leftarrow\,\Pi_{B_{*}}\Big{(}\boldsymbol{\alpha}_{i}-\eta(\boldsymbol{\alpha}_{i}-\boldsymbol{\alpha}_{i-1}+\mathbf{y}_{i}+\boldsymbol{\alpha}_{i}-\boldsymbol{\alpha}_{i+1}-\mathbf{y}_{i+1})\Big{)}~{}~{}.$
To underscore the advantage of the sub-gradient following method, let us briefly examine the condition number of the dual problem. Denote by $\mathsf{D}$ the difference matrix
$\mathsf{D}=\begin{bmatrix}-1&1&0&\dots&0&0\\ 0&-1&1&\dots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\dots&-1&1\end{bmatrix}\in\mathsf{R}^{n-1\,\times\,n}~{}~{}.$
@@
$\alpha_{i}=\frac{{(-1)}^{i}}{\sqrt{n-1}}~{}~{},$
we have $\|\mathsf{D}^{\top}\boldsymbol{\alpha}\|_{2}^{2}=\boldsymbol{\alpha}^{\top}\mathsf{D}\mathsf{D}^{\top}\boldsymbol{\alpha}=4-\frac{2}{n-1}$. Assume without loss of generality that $n$ is even and let us choose next
$\alpha_{i}=\frac{\frac{n}{2}-|i-\frac{n}{2}|}{\sqrt{n^{3}/12}}~{}~{},$
for which we have $\|\boldsymbol{\alpha}\|_{2}\geq 1$ and $\boldsymbol{\alpha}^{\top}\mathsf{D}\mathsf{D}^{\top}\boldsymbol{\alpha}=12/n^{2}$. Therefore, for $n\geq 3$ the largest eigenvalue of $\mathsf{D}\mathsf{D}^{\top}$ is at least $3$ whereas the smallest eigenvalue is at most $\mathcal{O}(n^{-2})$. Therefore, the condition number of the dual problem is $\Omega(n^{2})$ which renders dual ascent algorithms, and in particular DPG, slow to converge.
@@ Appendix C Upper Bounds on Slack Variables
Denote the optimum of the original problem by ${\mathbf{x}}_{1:n}^{*}$, we claim that $\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}\leq\|\mathbf{y}_{i+1}-\mathbf{y}_{i}\|_{\infty}$ and thus we can set $D_{i}=\|\mathbf{y}_{i+1}-\mathbf{y}_{i}\|_{\infty}$. If $\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}=0$, the bound clearly holds. Otherwise let us define,
$S=\{j:|x_{i+1,j}^{*}-x_{i,j}^{*}|=\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}\}~{}~{}.$
Consider $j\in S$ and let $s_{i,j}=\text{sign}(x_{i+1,j}^{*}-x_{i,j}^{*})$. From the optimality condition, there exists $\boldsymbol{\alpha}_{1:n-1}$ such that
@@
$=$ $y_{i+1,j}-y_{i,j}+\alpha_{i+1,j}-2\alpha_{i,j}+\alpha_{i-1,j}~{}~{}.$
@@
$\leq$ $|y_{i+1,j}-y_{i,j}|+|\alpha_{i+1,j}|+|\alpha_{i-1,j}|-2|\alpha_{i,j}|~{}~{}.$
@@
$\big{\|}{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\big{\|}_{\infty}$$\leq$ $\frac{1}{|S|}\Big{(}\sum_{j\in S}|y_{i+1,j}-y_{i,j}|+\|\boldsymbol{\alpha}_{i+1}\|_{1}+\|\boldsymbol{\alpha}_{i-1}\|_{1}-2\Big{)}$
$\leq$ $\|\mathbf{y}_{i+1}-\mathbf{y}_{i}\|_{\infty}~{}~{},$
where for obtaining the second inequality we used the fact that for $j\notin S:\alpha_{i,j}=0$ and $\|\boldsymbol{\alpha}_{i}\|_{1}=1$ since $\|{\mathbf{x}}_{i+1}^{*}-{\mathbf{x}}_{i}^{*}\|_{\infty}\neq 0$.
@@ Appendix D Iterative Algorithm for High-Order Variational Penalties
$\frac{1}{2}\|{\mathbf{x}}-\mathbf{y}\|^{2}+\sum_{i=1}^{n-2}\lambda_{i}\big{|}(x_{i+2}-x_{i+1})-(x_{i+1}-x_{i})\big{|}^{2}~{}~{}.$
@@
${\mathbf{x}}^{t+1}=\displaystyle\operatorname*{\arg\!\min}_{\mathbf{x}}\|{\mathbf{x}}-\mathbf{y}\|^{2}+\sum_{i=1}^{n-2}{\big{|}(x_{i+2}-x_{i+1})-(x_{i+1}-x_{i})\big{|}^{2}}\,/\,{\rho_{i}^{t}}~{}~{},$
where $\rho_{i}^{t}=\displaystyle\max\big{(}\big{|}(x_{i+2}^{t}-x_{i+1}^{t})-(x_{i+1}^{t}-x_{i}^{t})\big{|},\,\epsilon\big{)}$.
@@
$|(x_{i+2}-2x_{i+1}+x_{i})-(x^{\prime}_{i+2}-2x^{\prime}_{i+1}+x^{\prime}_{i})|^{2}\leq 4\left(|x_{i+2}-x^{\prime}_{i+2}|^{2}+2|x_{i+1}-x^{\prime}_{i+1}|^{2}+|x_{i}-x^{\prime}_{i}|^{2}\right)~{}~{}.$
@@
$\tilde{\mathsf{D}}=\begin{bmatrix}-1&1&0&\dots&0&0\\ 0&-1&1&\dots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\dots&-1&1\\ 0&0&0&\dots&0&0\\ \end{bmatrix}\in\mathsf{R}^{n\,\times\,n}~{}~{}.$
@@
$\operatorname*{\arg\!\min}_{\mathbf{x}}\frac{1}{2}\big{\|}{\mathbf{x}}-\mathbf{y}\big{\|}^{2}+\big{\|}\big{[}\mathsf{\tilde{D}}^{k}{\mathbf{x}}\big{]}_{1:n-k}\big{\|}_{1}~{}~{}.$
In terms of run time, we “pay” a factor of $4^{k}$ for the number of iterations required to converge and each iteration takes $\mathcal{O}(k^{2}n)$ for solving the linear system. In comparison, when using the dual proximal gradient method, the iteration complexity grows as $\mathcal{O}(\gamma^{k})$ where $\gamma=\Omega(n^{2})$ is the condition number of the first order problem. This implies a much inferior time complexity of $\mathcal{O}(n^{2k})$ for DPG. Pseudocode for $k>0$ is provided below.
Algorithm 2 $k^{\text{th}}$-Order Untangled Lariat